# Hamiltonian stochastic Lie systems and applications

### v2

@@ 5 Stability for stochastic Hamiltonian systems
is the associated Stratonovich operator. As in previous sections, it is assumed that initial conditions are values of $M$ which are taken with probability equal to one. Note that coefficients of the Stratonovich operator are to depend only on $t$ and $\Gamma$. Stochastic differential equations of this type are common in the literature [11].
@@
In this case, the stochastic disturbance, which is described by $\mathfrak{S}_{\mathcal{B}}$, does not act at the equilibrium point $\Gamma_{e}$. Note that if $\mathfrak{S}_{\mathcal{B}}$ is not assumed to be zero at the equilibrium point for every $t\in\mathbb{R}$, solutions may move away from the point depending on the values of the stochastic variable (see Figure 2).
@@
Let us provide a relevant analogue of the method used in Hamiltonian symplectic systems to verify stability [1]. But before that, let us provide a definition for strongly conserved quantities for Hamiltonian stochastic differential equations (see [40, Definition 2.2]).
@@
Note that strongly conserved quantities can also be defined for systems (5.2) whose Stratonovich operator depend also on $B$. The most relevant result for our purposes in the study of stability of Hamiltonian stochastic Lie systems is the following proposition (see [40, Theorem 2.15] and references therein).
@@
(Stochastic Dirichlet’s Criterion) Assume that there exists a function $f\in\mathscr{C}^{\infty}(M)$ so that $\mathrm{d}f_{\Gamma_{e}}=0$ and that the quadratic form $\mathrm{d}^{2}f_{\Gamma_{e}}$ is positive or negative definite. If $f$ is a strongly conserved quantity on the solutions of (5.2), then the equilibrium point $\Gamma_{e}$ is almost surely stable.
@@
where $B\colon\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}^{\ell}$ is a semi-martingale and $\mathfrak{H}:\mathrm{T}\mathbb{R}^{\ell}\times M\rightarrow\mathrm{T}M$ is a Stratonovich operator such that
@@
where $Y_{1},\ldots,Y_{\ell}$ stand for $t$-dependent Hamiltonian vector fields relative to a symplectic form $\omega$ on $M$ with $t$-dependent Hamiltonian functions $\widetilde{h}_{1},\ldots,\widetilde{h}_{\ell}\in\mathscr{C}^{\infty}(\mathbb{R}\times M)$, respectively. The symplectic structure in this case allows for the study of the system via the Hamiltonian functions of the system, which gives powerful methods to study their properties. Note that (5.3) has indeed an associated function $\widetilde{h}\colon\mathbb{R}\times M\rightarrow\mathbb{R}^{\ell}$ of the form
$\widetilde{h}=(\widetilde{h}_{1},\ldots,\widetilde{h}_{\ell}).$
Recall that $\widetilde{h}_{1}$ is normally called the Hamiltonian of the system, but Hamiltonian stochastic differential equations have several associated Hamiltonians. In fact, note that $\widetilde{h}_{1}$ need not be conserved even when it is $t$-independent. It is immediate that it is not a constant of the motion in general as $\{\widetilde{h}_{1},\widetilde{h}_{i}\}$ not need to be zero in general. Notwithstanding, it is remarkable that a $f\in\mathscr{C}^{\infty}(M)$ is a strong constant of motion for (4.5) if
@@
Let us assume that there exists a function $f$ that is a constant of motion of $\delta\Gamma$ and it also satisfies that it has a minimum at an equilibrium point. Then, it is immediate that $\delta\Gamma$ has a stable equilibrium point.
It is simple to apply the above results to linear or affine stochastic differential equations depending on a semi-martingale $B$, which are stochastic Lie systems, and to choose cases with a Vessiot–Guldberg Lie algebra of Hamiltonian vector fields relative to a symplectic structure.
@@ 6 Relative equilibrium points and stochastic Hamiltonian Lie systems
In short, relative equilibrium points for a system of differential equations on a manifold $M$, in general, and for a system of stochastic differential ones, in particular, are points where the dynamics is generated by symmetries of the differential equations. In this section, we define relative equilibrium points for Hamiltonian stochastic differential equations in a Stratonovich form of the particular type (5.3) using the theory in [38, 40].
Let us start by reviewing the notion of symmetries for stochastic differential equations and symplectic reduction for Hamiltonian stochastic differential equations.
@@
More generally, a Lie group action $\Phi:G\times M\rightarrow M$ is a symmetry of the stochastic differential equation induced by $\mathfrak{S}$ if $\Phi_{g}:\Gamma\in M\mapsto\Phi(g,\Gamma)\in M$ is a symmetry of $\mathfrak{S}$ for every $g\in G$. Similarly, we say that $\phi$ and $\Phi$ are symmetries of the Stratonovich operator $\mathfrak{S}$.
@@
Note the slight change of notation for the Stratonovich operator in Definition 6.1, where the first entry is an element of $\mathrm{T}\mathbb{R}^{\ell}$ instead of an element of $\mathbb{R}^{\ell}$ as in previous sections.
The relevance of symmetries of Stratonovich operators is due to the fact that they transform a particular solution of the stochastic differential equations related to it to another particular solution of the same stochastic differential equation [38, Theorem 2.2].
Let us give a $t$-dependent generalisation of the stochastic Marsden–Meyer–Weinstein reduction for Hamiltonian stochastic systems as introduced in [41, Theorem 3.1] for the case of Hamiltonian stochastic systems (5.3), which essentially amounts to applying a standard reduction for each particular $t\in\mathbb{R}$.
Let $(M,\omega)$ be a symplectic manifold and let $\Phi\colon G\times M\rightarrow M$ be a Lie group action admitting a coadjoint equivariant momentum map $J\colon M\rightarrow\mathfrak{g}^{*}$ being also a Lie group symmetry of the Hamiltonian functions $\widetilde{h}\colon\mathbb{R}\times M\rightarrow\mathbb{R}^{\ell}$ of a Stratonovich operator $\mathfrak{H}$. Every regular $\mu\in\mathfrak{g}^{*}$ of $J$ gives rise to a function $h_{\mu}\colon\mathbb{R}\times J^{-1}(\mu)/G_{\mu}\rightarrow\mathbb{R}^{\ell}$ on the manifold $M_{\mu}=J^{-1}(\mu)/G_{\mu}$ determined by the equality $h_{\mu}(t,\pi_{\mu}(\cdot))=h(t,\iota_{\mu}(\cdot))$, where $\iota_{\mu}\colon J^{-1}(\mu)\hookrightarrow M$ is the natural immersion and we assume $G_{\mu}$ to act freely and properly on $J^{-1}(\mu)$. In turn, this induces a stochastic Hamiltonian system on the symplectic reduced space $(M_{\mu}:=J^{-1}(\mu)/G_{\mu},\omega_{\mu})$ whose Stratonovich operator $\mathfrak{H}_{\mu}\colon\mathrm{T}\mathbb{R}^{\ell}\times M_{\mu}\to\mathrm{T}M_{\mu}$ is given by (c.f. [38, Theorem 3.1])
@@
Note that $\mathrm{T}_{\Gamma}\pi_{\mu}$ is assumed to act on each component of $\mathfrak{H}(t,\Gamma)\colon\mathrm{T}_{t}\mathbb{R}^{\ell}\rightarrow\mathrm{T}_{\Gamma}M$ for each $t\in\mathbb{R}$ in the natural manner. Moreover, if $\Gamma$ is a solution of the stochastic Hamiltonian system associated with $\mathfrak{H}$ with initial condition $\Gamma_{0}\subset J^{-1}(\mu)$ at $t=0$, then so is $\Gamma_{\mu}:=\pi_{\mu}(\Gamma)$ with respect to $\mathfrak{H}_{\mu}$, with initial condition $\pi_{\mu}(\Gamma_{0})$ at $t=0$.
@@
$\mathfrak{H}_{\mu}(t,[\Gamma]))=\left(\sum_{\alpha=1}^{r}b^{1}_{\alpha}(t)X_{h^{\mu}_{\alpha}}([\Gamma]),\ldots,\sum_{\alpha=1}^{r}b^{\ell}_{\alpha}(t)X_{h^{\mu}_{\alpha}}([\Gamma])\right),$
where the $h^{\mu}_{\alpha}\in\mathscr{C}^{\infty}(M_{\mu})$ are determined by the conditions $h^{\mu}_{\alpha}\circ\pi_{\mu}=h_{\alpha}\circ\iota_{\mu}$ from the generators of the Lie algebra of Hamiltonian functions $h_{1},\ldots,h_{r}$ of the initial Hamiltonian stochastic Lie system. Note that in the case of Hamiltonian stochastic Lie systems with a Vessiot–Guldberg Lie algebra $V$, we require all related Hamiltonian functions, $h_{1},\ldots,h_{r}$ to be invariant relative to the action of the Lie group $\Phi\colon G\times M\rightarrow M$.
The above discussion suggests us the following definition.
@@
$\mathfrak{H}_{i}(t,\Gamma_{\rm rel})\in D_{\Gamma_{\rm rel}},\qquad i=1,\ldots,\ell\,,\qquad\forall t\in\mathbb{R}\,,$ (6.1)
where $D$ is the distribution generated by the fundamental vector fields of a Lie group action of symmetries $\Phi\colon G\times M\rightarrow M$ with a momentum map $J\colon M\rightarrow\mathfrak{g}^{*}$.
@@
Let $\mu=J(\Gamma_{\rm rel})$ in the above definition. If the Hamiltonian stochastic differential equation (5.3) can be reduced, then $\Gamma_{\rm rel}\in M$ is contained in $J^{-1}(\mu)$. If it is assumed to be a submanifold of $M$, the evolution of $\mathfrak{H}$ is contained within $J^{-1}(\mu)$. This implies that the right-hand side of (6.1) belongs indeed to $D_{\Gamma_{\rm rel}}\cap\mathrm{T}_{\Gamma_{\rm rel}}J^{-1}(\mu)$ Moreover, the restriction of $\mathfrak{H}$ to it can be reduced to $J^{-1}(\mu)/G_{\mu}$. Moreover, the $\mathfrak{H}_{\mu}(t,\pi_{\mu}(J^{-1}(\mu)))=0$. In other words, one has the following proposition, whose inverse is analogue to its deterministic case.
@@
Let $\Gamma_{\rm rel}\in M$ be a relative equilibrium point for a stochastic Hamiltonian system (5.3) and let $J(\Gamma_{\rm rel})=\mu$ be a regular point of $J$, let $G_{\mu}$ act on $J^{-1}(\mu)$ freely and properly so that $J^{-1}(\mu)/G_{\mu}$ becomes a manifold. Then, the reduction of (5.3) to $M_{\mu}=J^{-1}(\mu)/G_{\mu}$ is such that $\pi_{\mu}(\Gamma_{\rm rel})$ is an equilibrium point of the reduced system and vice versa.
@@
Let us prove the direct part. At the critical points $(\xi_{t},\Gamma_{\rm rel})$, the functions $\widehat{h}_{t}(\xi_{t},\cdot):M\rightarrow\mathbb{R}^{\ell}$ have a critical point $\Gamma_{\rm rel}\in M$. Hence, each coordinate of $\widehat{h}_{t}(\xi_{t},\cdot)$ satisfies that $\mathrm{d}(\widetilde{h}_{t}^{\alpha}-\langle J(\cdot),\xi^{\alpha}_{t}\rangle)|_{\Gamma_{\rm rel}}=0$. This implies that $(X_{\widetilde{h}_{t}^{\alpha}})|_{\Gamma_{\rm rel}}=X_{J_{\xi^{\alpha}_{t}}}|_{\Gamma_{\rm rel}}$ and the point $\Gamma_{\rm rel}$ is a relative equilibrium point of $\mathfrak{H}$. Conversely, if $\Gamma_{\rm rel}$ is a relative equilibrium point to $\mathfrak{H}$, at each coordinate of $\mathfrak{H}$, one has that $(X_{\widetilde{h}^{\alpha}})_{t}(\Gamma_{\rm rel})=(\xi_{t})_{M}^{\alpha}(\Gamma_{\rm rel})$ for certain $\xi_{t}^{1},\ldots,\xi_{t}^{\ell}\in\mathfrak{g}$ with $t\in\mathbb{R}$. Consequently, one has that $X_{\widetilde{h}_{t}^{\alpha}-\langle J,\xi^{\alpha}_{t}\rangle}=0$. This implies that each one of these vector fields has a Hamiltonian function given by $\widetilde{h}_{t}^{\alpha}-\langle J-\mu_{e},\xi_{t}^{\alpha}\rangle$ which has a critical point at the given point $\Gamma_{\rm rel}$. ∎

### v3

@@ 5 Stability for stochastic Hamiltonian systems
is the associated Stratonovich operator. As in previous sections, it is assumed that initial conditions are values in $M$ chosen with probability equal to one. Note that coefficients of the Stratonovich operator are considered to depend only on $t$ and $\Gamma$. Stochastic differential equations of this type are common in the literature [15].
@@
In this case, the stochastic disturbance, which is described by $\mathfrak{S}_{\mathcal{B}}$, does not act at the equilibrium point $\Gamma_{e}$. Note that if $\mathfrak{S}_{\mathcal{B}}$ is not assumed to be zero at the equilibrium point for every $t\in\mathbb{R}$, solutions may move away from that point depending on the values of the stochastic variable (see Figure 2).
@@
Let us provide a relevant analogue of the method used in Hamiltonian symplectic systems to verify stability [1]. Before that, let us provide a definition of strongly conserved quantities for Hamiltonian stochastic differential equations (see [46, Definition 2.2]).
@@
Strongly conserved quantities can also be defined for systems (5.2) whose Stratonovich operator also depend on $B$. The most relevant result for our purposes in the study of stability of Hamiltonian stochastic Lie systems is the following proposition (see [46, Theorem 2.15] and references therein).
@@
(Stochastic Dirichlet’s Criterion) Assume that there exists a function $f\in\mathscr{C}^{\infty}(M)$ such that $\mathrm{d}f_{\Gamma_{e}}=0$ and that the quadratic form $\mathrm{d}^{2}f_{\Gamma_{e}}$ is positive or negative definite. If $f$ is a strongly conserved quantity on the solutions of (5.2), then the equilibrium point $\Gamma_{e}$ is almost surely stable.
@@
where $B\colon\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}^{\ell}$ is a semi-martingale and $\mathfrak{H}(t,\Gamma):\mathrm{T}_{B}\mathbb{R}^{\ell}\rightarrow\mathrm{T}_{\Gamma}M$ is a Stratonovich operator such that
@@
where $Y_{1},\ldots,Y_{\ell}$ stand for $t$-dependent Hamiltonian vector fields relative to a symplectic form $\omega$ on $M$ with $t$-dependent Hamiltonian functions $\widetilde{h}_{1},\ldots,\widetilde{h}_{\ell}\in\mathscr{C}^{\infty}(\mathbb{R}\times M)$, respectively. The symplectic structure in this case allows the study of the system via its Hamiltonian functions, which provides powerful methods to study their properties. Note that (5.3) has indeed an associated function $\widetilde{h}\colon\mathbb{R}\times M\rightarrow\mathbb{R}^{\ell}$ of the form
[…]
Recall that $\widetilde{h}_{1}$ is normally called the Hamiltonian of the system, but Hamiltonian stochastic differential equations have several associated Hamiltonians. In fact, $\widetilde{h}_{1}$ need not be conserved even when it is $t$-independent. It is immediate that it is not a constant of motion in general as $\{\widetilde{h}_{1},\ldots,\widetilde{h}_{\ell}\}$ not need to be zero in general. Notwithstanding, it is remarkable that an $f\in\mathscr{C}^{\infty}(M)$ is a strong constant of motion for (4.7) if (cf. [46])
@@
Let us assume that there exists a function $f$ that is a constant of motion of $\delta\Gamma$ and it also satisfies that it has a strict minimum at an equilibrium point. Then, it is immediate that $\delta\Gamma$ has a stable equilibrium point.
It is immediate to apply the above results to linear or affine stochastic differential equations, which are stochastic Lie systems, and to choose cases with a Vessiot–Guldberg Lie algebra of Hamiltonian vector fields relative to a symplectic structure.
Let us consider a particular case of the the stochastic differential equation on $\mathbb{R}^{n}$ with a Wiener process $W_{1}$ given in (4.4) for $2a(t)=a^{\prime}(t)=2$, $2b(t)=b^{\prime}(t)=-2\omega^{2}$ and $n=1$. Then, (4.4) can now be related to a Vessiot–Guldberg Lie algebra $V$ on $\mathrm{T}^{*}\mathbb{R}$ spanned by
$X_{1}=y\frac{\partial}{\partial x}\,,\qquad X_{2}=\frac{1}{2}\left(x\frac{\partial}{\partial x}-y\frac{\partial}{\partial y}\right)\,,\qquad X_{3}=-x\frac{\partial}{\partial y}$
spanning a Lie algebra isomorphic to $\mathfrak{sl}_{2}$ of Hamiltonian vector fields relative to the symplectic form $\omega=\mathrm{d}x\wedge\mathrm{d}y\,.$ Hence, $X_{1},X_{2},X_{3}$ have Hamiltonian functions given by
$h_{1}=\frac{y^{2}}{2}\,,\qquad h_{2}=\frac{1}{2}\ xy\,,\qquad h_{3}=\frac{x^{2}}{2}\,.$
relative to $\omega$ that span a Lie algebra isomorphic to $\mathfrak{sl}_{2}$. Now, (4.4) can be related to Stratonovich operator of the form
$\mathfrak{H}(x,y,t,W)=\left(X_{1}+\omega^{2}X_{3},2X_{1}+2\omega^{2}X_{3}\right),$
which turns (4.4) into a Hamiltonian stochastic Lie system and $V$ into an associated Vessiot–Guldberg Lie algebra isomorphic to $\mathfrak{sl}_{2}$. Note that this system has two Hamiltonians
$\widetilde{h}=(y^{2}/2+\omega^{2}x^{2}/2,y^{2}+\omega^{2}x^{2}).$
Then, $y^{2}/2+\omega^{2}/2$ is the Hamiltonian of the system. The point $(0,0)$ is an equilibrium point and the function $f=y^{2}/2+\omega^{2}x^{2}/2$ is strongly conserved because $\{f,f\}=\{2f,f\}=0$ relative to the Poisson bracket induced by $\omega$. Since $d^{2}f$ is positive definite, then $(0,0)$ is almost surely stable.
@@ 6 Relative equilibrium points and stochastic Hamiltonian Lie systems
In short, relative equilibrium points for a system of differential equations on a manifold $M$, in general, and for a system of stochastic differential equations, in particular, are points where the dynamics is generated by Lie group actions such that the elements of the group are understood as symmetries of the differential equations. In this section, we define for the first time relative equilibrium points for Hamiltonian stochastic differential equations in Stratonovich form of the particular type (5.3) using the theory in [46, 47].
Let us start by reviewing the notion of symmetries for stochastic differential equations and symplectic reduction for Hamiltonian stochastic differential equations (see [47, Section 2] for details).
@@
More generally, a Lie group action $\Phi:G\times M\rightarrow M$ is a Lie group of symmetries of the stochastic differential equation induced by $\mathfrak{S}$ if $\Phi_{g}:\Gamma\in M\mapsto\Phi(g,\Gamma)\in M$ is a symmetry of $\mathfrak{S}$ for every $g\in G$. Similarly, we say that $\phi$ (resp. $\Phi$) are symmetries (resp. a Lie group of symmetries) of the Stratonovich operator $\mathfrak{S}$.
@@
Note the slight change of notation for the Stratonovich operator in Definition 6.1, where for instance the first entry is an element of $\mathrm{T}\mathbb{R}^{\ell}$ instead of an element of $\mathbb{R}^{\ell}$ as in previous sections. It is worth stressing that the term ‘Lie group symmetries’ also appears in the literature on standard and stochastic PDEs with the same meaning of ‘Lie group of symmetries’ (cf. [6]).
The relevance of the symmetries of Stratonovich operators is due to the fact that they transform a particular solution of the stochastic differential equations related to it to another particular solution of the same stochastic differential equation [47, Theorem 2.2]. In the case of a Lie group of symmetries of a stochastic differential equation induced by $\mathfrak{S}$, then every element of the associated Lie group can be understood as a symmetry of the stochastic differential equation.
Let us give a $t$-dependent generalisation of the stochastic Marsden–Meyer–Weinstein reduction for Hamiltonian stochastic systems as introduced in [47, Theorem 3.1] for the case of Hamiltonian stochastic systems (5.3), which essentially amounts to applying a standard reduction for each particular $t\in\mathbb{R}$.
Let $(M,\omega)$ be a symplectic manifold and let $\Phi\colon G\times M\rightarrow M$ be a Lie group action admitting a coadjoint equivariant momentum map $J\colon M\rightarrow\mathfrak{g}^{*}$ being also a Lie group of symmetries of the Hamiltonian functions $\widetilde{h}\colon\mathbb{R}\times M\rightarrow\mathbb{R}^{\ell}$ of a Stratonovich operator $\mathfrak{H}$. Every regular $\mu\in\mathfrak{g}^{*}$ of $J$ gives rise to a function $h_{\mu}\colon\mathbb{R}\times J^{-1}(\mu)/G_{\mu}\rightarrow\mathbb{R}^{\ell}$ on the manifold $M_{\mu}=J^{-1}(\mu)/G_{\mu}$ determined by the equality $h_{\mu}(t,\pi_{\mu}(\cdot))=h(t,\iota_{\mu}(\cdot))$ for every $t\in\mathbb{R}$, where $\iota_{\mu}\colon J^{-1}(\mu)\hookrightarrow M$ is the natural immersion, we assume the isotropy subgroup $G_{\mu}$ of the coadjoint action at $\mu$ to act freely and properly on $J^{-1}(\mu)$, and $\pi_{\mu}:J^{-1}(\mu)\rightarrow M_{\mu}$ is the natural projection. Moreover, one obtains a unique symplectic form $\omega_{\mu}$ on $M_{\mu}$ induced by the relation $\pi_{\mu}^{*}\omega_{\mu}=\iota_{\mu}^{*}\omega$. In turn, this induces a stochastic Hamiltonian system on the symplectic reduced space $(M_{\mu},\omega_{\mu})$ whose Stratonovich operator $\mathfrak{H}_{\mu}\colon\mathrm{T}\mathbb{R}^{\ell}\times M_{\mu}\to\mathrm{T}M_{\mu}$ is given by (cf. [47, Theorem 3.1])
@@
Moreover, if $\Gamma$ is a solution of the stochastic Hamiltonian system associated with $\mathfrak{H}$ with initial condition $\Gamma_{0}\subset J^{-1}(\mu)$ at $t=0$, then so is $\Gamma_{\mu}:=\pi_{\mu}(\Gamma)$ with respect to $\mathfrak{H}_{\mu}$, with initial condition $\pi_{\mu}(\Gamma_{0})$ at $t=0$.
@@
$\mathfrak{H}_{\mu}(t,[\Gamma]))=\left(\sum_{\alpha=1}^{r}b_{1}^{\alpha}(t)X_{h^{\mu}_{\alpha}}([\Gamma]),\ldots,\sum_{\alpha=1}^{r}b_{\ell}^{\alpha}(t)X_{h^{\mu}_{\alpha}}([\Gamma])\right),$
where the $h^{\mu}_{\alpha}\in\mathscr{C}^{\infty}(M_{\mu})$ are determined by the conditions $h^{\mu}_{\alpha}\circ\pi_{\mu}=h_{\alpha}\circ\iota_{\mu}$ from the generators of the Lie–Hamiltonian Lie algebra of functions $h_{1},\ldots,h_{r}$ of the initial Hamiltonian stochastic Lie system. Note that in the case of Hamiltonian stochastic Lie systems with a Vessiot–Guldberg Lie algebra $V$, we require all related Hamiltonian functions, $h_{1},\ldots,h_{r}$ to be invariant relative to the action of the Lie group $\Phi\colon G\times M\rightarrow M$. This condition is always satisfied for the smallest Lie algebra $V^{X}$ as it follows from the fact that $\mathfrak{H}$ can be reduced for each time $t$.
The above discussion suggests the following definition.
@@
$\mathfrak{H}_{\alpha}(t,\Gamma_{\rm rel})\in D_{\Gamma_{\rm rel}},\qquad\alpha=1,\ldots,\ell\,,\qquad\forall t\in\mathbb{R}\,$, (6.1)
where $D$ is the distribution generated by the fundamental vector fields of a Lie group action of symmetries $\Phi\colon G\times M\rightarrow M$ of (5.3) with a momentum map $J\colon M\rightarrow\mathfrak{g}^{*}$.
@@
Let $\mu=J(\Gamma_{\rm rel})$ in the above definition. If the Hamiltonian stochastic differential equation (5.3) can be reduced, then $\Gamma_{\rm rel}\in M$ is contained in $J^{-1}(\mu)$. If $J^{-1}(\mu)$ is a submanifold of $M$, the evolution of $\mathfrak{H}$ is contained within $J^{-1}(\mu)$. This implies that the right-hand side of (6.1) belongs indeed to $D_{\Gamma_{\rm rel}}\cap\mathrm{T}_{\Gamma_{\rm rel}}J^{-1}(\mu)$. Moreover, the restriction of $\mathfrak{H}$ to it can be reduced to $J^{-1}(\mu)/G_{\mu}$ and the $\mathfrak{H}_{\mu}(t,\pi_{\mu}(J^{-1}(\mu)))=0$. In other words, one has the following proposition, whose inverse is analogue to its deterministic case [27].
@@
Let $\Gamma_{\rm rel}\in M$ be a relative equilibrium point for a stochastic Hamiltonian system (5.3) and let $J(\Gamma_{\rm rel})=\mu$ be a regular point of $J$. Let $G_{\mu}$ act on $J^{-1}(\mu)$ freely and properly so that $J^{-1}(\mu)/G_{\mu}$ becomes a manifold. Then, the reduction of (5.3) to $M_{\mu}=J^{-1}(\mu)/G_{\mu}$ is such that $\pi_{\mu}(\Gamma_{\rm rel})$ is an equilibrium point of the reduced system and vice versa.
@@
Let us prove the direct part. At the points $(\xi_{t},\Gamma_{\rm rel})$, the coordinates of the functions $\widehat{h}_{t}(\xi_{t},\cdot):M\rightarrow\mathbb{R}^{\ell}$ have a critical point $\Gamma_{\rm rel}\in M$. Hence, each coordinate of $\widehat{h}_{t}(\xi_{t},\cdot)$ satisfies that $\mathrm{d}(\widetilde{h}_{t}^{\alpha}-\langle J(\cdot),\xi^{\alpha}_{t}\rangle)|_{\Gamma_{\rm rel}}=0$. This implies that $(X_{\widetilde{h}_{t}^{\alpha}})({\Gamma_{\rm rel})=X_{J_{\xi^{\alpha}_{t}}}}(\Gamma_{\rm rel})$ and the point $\Gamma_{\rm rel}$ is a relative equilibrium point of $\mathfrak{H}$. Conversely, if $\Gamma_{\rm rel}$ is a relative equilibrium point to $\mathfrak{H}$, at each coordinate of $\mathfrak{H}$, one has that $(X_{\widetilde{h}^{\alpha}})_{t}(\Gamma_{\rm rel})=(\xi_{t})_{M}^{\alpha}(\Gamma_{\rm rel})$ for certain $\xi_{t}^{1},\ldots,\xi_{t}^{\ell}\in\mathfrak{g}$ with $t\in\mathbb{R}$. Consequently, one has that $X_{\widetilde{h}_{t}^{\alpha}-\langle J,\xi^{\alpha}_{t}\rangle}=0$. This implies that each one of these vector fields has a Hamiltonian function given by $\widetilde{h}_{t}^{\alpha}-\langle J-\mu_{e},\xi_{t}^{\alpha}\rangle$ which has a critical point at the given point $\Gamma_{\rm rel}$. ∎
Let us consider a particular case of the the stochastic differential equation on $\mathbb{R}^{n}$ with a Wiener process $W_{1}$ given in (4.4) for $2a(t)=a^{\prime}(t)=2$, $2b(t)=b^{\prime}(t)=-2\omega^{2}$, any $n$, and a constant $\omega$ (see [25] for similar models). Then, (4.4) can now be related to a Vessiot–Guldberg Lie algebra $V$ on $\mathrm{T}^{*}\mathbb{R}^{n}$ spanned by the vector fields (4.5) isomorphic to $\mathfrak{sl}_{2}$ and consisting of Hamiltonian vector fields relative to the symplectic form $\omega=\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}\,.$ Hence, $X_{1},X_{2},X_{3}$ have Hamiltonian functions given by (4.6) relative to $\omega$ that span a Lie algebra of functions isomorphic to $\mathfrak{sl}_{2}$. Now, (4.4) can be related to Stratonovich operator of the form
$\mathfrak{H}(x,y,t,W)=\left(X_{1}+\omega^{2}X_{3},2X_{1}+2\omega^{2}X_{3}\right),$
which turns (4.4) into a Hamiltonian stochastic Lie system and $V$ into an associated Vessiot–Guldberg Lie algebra. Note that this system has two Hamiltonians
$\widetilde{h}=(h_{1},h_{2})=\left(\sum_{i=1}^{n}(y_{i}^{2}/2+\omega^{2}x_{i}^{2}/2),\sum_{i=1}^{n}(y_{i}^{2}+\omega^{2}x_{i}^{2})\right).$
Let us consider the Lie symmetries $Y_{k}=-\omega^{2}x_{k}\frac{\partial}{\partial y_{k}}+y_{k}\frac{\partial}{\partial x_{k}}$ for $k=1,\ldots,s$ of the above Hamiltonian system. In other words, $Y_{1},\ldots,Y_{s}$ are Hamiltonian relative to $\omega$ and are Lie symmetries of $\mathfrak{H}$. Moreover, $Y_{1},\ldots,Y_{s}$ commute between themselves giving rise to an Abelian Lie algebra of vector fields whose integration gives rise to a Lie group action on $\mathrm{T}\mathbb{R}^{n}$ and an associated momentum map
${\rm J}:(x_{1},y_{1},\ldots,x_{n},y_{n})\in\mathrm{T}\mathbb{R}^{n}\mapsto(x_{1}^{2}/2+\omega^{2}y_{1}^{2}/2,\ldots,x_{s}^{2}/2+\omega^{2}y_{s}^{2}/2)\in\mathbb{R}^{s*}.$
Since $Y_{1},\ldots,Y_{s}$ are Hamiltonian Lie symmetries of $(h_{1},h_{2})$, the system can be reduced. Following the approach in [47], one projects the Hamiltonian stochastic system on $J^{-1}(c_{1},\ldots,c_{s})$ with $\prod_{i=1}^{s}c_{i}\neq 0$. The latter condition is employed to ensure that $(c_{1},\ldots,c_{s})$ is a regular value of $J$ and $J^{-1}(c_{1},\ldots,c_{s})$ is a submanifold. Note that $Y_{1}\wedge\ldots\wedge Y_{s}\neq 0$ and $Y_{1},\ldots,Y_{s}$ are tangent to $J^{-1}(c_{1},\ldots,c_{s})$, which gives rise to a reduced stochastic system on $J^{-1}(c_{1},\ldots,c_{s})/\mathbb{R}^{s}\simeq\mathrm{T}\mathbb{R}^{n-s}$, which is Hamiltonian relative to a symplectic form $\omega=\sum_{i=s+1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}$. More exactly, the reduced stochastic system reads
$\delta\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}=\begin{pmatrix}0&1\\ -\omega^{2}&0\end{pmatrix}\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}\delta t+\begin{pmatrix}0&2\\ -2\omega^{2}&0\end{pmatrix}\begin{pmatrix}x_{i}\\ y_{i}\end{pmatrix}\circ\delta W_{1}\,,\qquad i=s+1,\ldots,n\,.$ (6.2)
Note that the equilibrium points of the above system are those points such that $x_{s+1}=y_{s+1}=\ldots=x_{n}=y_{n}=0$, which are the projections of the original ones of the form
$(x_{1},y_{1},\ldots,x_{s},y_{s},\stackrel{{\scriptstyle(n-s)-{\rm pairs}}}{{\overbrace{0,0,\ldots,0,0}}}).$
Note that these points are indeed the relative equilibrium points of our initial stochastic system, where the components of $\mathfrak{H}$ can be written as a linear combination of $Y_{1},\ldots,Y_{s}$. Indeed, at these points $X_{1}+\omega^{2}X_{3}=\sum_{i=1}^{s}Y_{i}$ and $2X_{1}+2\omega^{2}X_{3}=\sum_{i=1}^{s}2Y_{i}$. As stated in Proposition 6.5, one has that at these points $h_{1}-\sum_{i=1}^{s}x_{i}^{2}/2+\omega^{2}y_{i}^{2}/2$ and $h_{2}-\sum_{i=1}^{s}x_{i}^{2}+\omega^{2}y_{i}^{2}$ have equilibrium points.